OpenStudy (anonymous):

How do you convert y + 13=4(x+2) into slope intercept form?

OpenStudy (anonymous):

Simplify the equation first...use the distributive property on the right side.

OpenStudy (anonymous):

y+13 =4x+8 Slope intercept form is y=mx+b. You must get y by itself by moving 13 to the other side of the equation and simplifying.

OpenStudy (anonymous):

okay so would do 4x-13?

OpenStudy (anonymous):

not quite...you'd have y+13-13 = 4x +8-13 when you simplify, you'd get y=4x-5

OpenStudy (anonymous):

You lost the +8...don't throw it away, just combine it with the -13

OpenStudy (anonymous):

why did the 13 turn negative?

OpenStudy (nincompoop):

4 + 6 = 10 6 = 10 - 4 see how 4 turned negative?

OpenStudy (anonymous):

You had to get y by itself. y+13 = 4x+8 y+13-13 = 4x+8-13 y+0 = 4x-5 y=4x-5 does that make sense

OpenStudy (nincompoop):

one invisible step that we don't normally show in simple arithmetic is this part 4 - 4 + 6 = 10 - 4

OpenStudy (nincompoop):

to move any variable or value from one side to another, its property changes into opposite from + to - from - to + from x to ÷ from ÷ to x

OpenStudy (anonymous):

yes, it makes a lot more sense now, thank you dss and nincompoop i did not know that helps a lot! thank you to you both. :)

OpenStudy (anonymous):

You're welcome. Good questions